Find the center and the radius of x^2+y^2+2 x+6 y+1=0
Answer by lwsshak3(11628)

Find the center and the radius
x^2+y^2+2 x+6 y+1=0
complete the square
(x^2+2x+1)+(y^2+6y+9)=-1+1+9
(x+1)^2+(y+3)^2=9
Center: (-1, -3)
Radius=√9=3